Epsilon Airlines services predominantly the eastern and southeastern united States. The vast majority of Epsilon’s customers make reservations through Epsilon’s website, but a small percentage of customers make reservations via phones. Epsilon employs call center personnel to handle these reservations and to deal with website reservation system problems and for the rebooking of flights for customers whose plans have changed or whose travel is disrupted. Staffing the call center appropriately is a challenge for Epsilon’s management team. Having too many employees on hand is a waste of money, but having too few results in very poor customer service and the potential loss of customers.

Epsilon analysts have estimated the minimum number of call center employees needed by day of the week for the upcoming vacation season (June, July, and the first two weeks of August). These estimates are as follows:


Day
Minimum Number of
Employees Needed
Monday 50
Tuesday 90
Wednesday 45
Thursday 55
Friday 75
Saturday 55
Sunday 45

The call center employees work for five consecutive days and then have two consecutive days off. An employee may start work on any day of the week. Each call center employee receives the same salary. Assume that the schedule cycles and ignore start up and stopping of the schedule.

Develop a model that will minimize the total number of call center employees needed to meet the minimum requirements.

Let Xi = the number of call center employees who start work on day i
      (i = 1 = Monday, i = 2 = Tuesday...)
Min X1 + X2 + X3 + X4 + X5 + X6 + X7
s.t.
X1 + X4+ X5+ X6+ X7 ≤, ≥, or = ________
X1 + X2+ X5+ X6+ X7 ≤, ≥, or = ________
X1 + X2+ X3+ X6+ X7 ≤, ≥, or = ________
X1 + X2+ X3+ X4+ X7 ≤, ≥, or = ________
X1 + X2+ X3+ X4+ X5 ≤, ≥, or = ________
X2 + X3+ X4+ X5+ X6 ≤, ≥, or = ________
X3 + X4+ X5+ X6+ X7 ≤, ≥, or = ________
X1, X2, X3, X4, X5, X6, X7 0

Find the optimal solution.

X1 = ________
X2 = ________
X3 = ________
X4 = ________
X5 = ________
X6 = ________
X7 = ________

Total Number of Employees = ________

Give the number of call center employees that exceed the minimum required.

Excess employees:

Monday = ________
Tuesday = ________
Wednesday = ________
Thursday = ________
Friday = ________
Saturday = ________
Sunday = ________
Answer #1

Here the objective function will be x1+x2+x3+x4+x5+x6+x7 and this has to be minimized.

A person joining on Thursday will be available till Monday. So the constraint will be:

X1+x4+x5+x6+x7>=50

Similarly other constraints will be:

X1+x2+x5+x6+x7>=90

X1+x2+x3+x6+x7>=45

X1+x2+x3+x4+x7>=55

X1+x2+x3+x4+x5>=75

x2+x3+x4+x5+x6>=55

x3+x4+x5+x6+x7>=45

Lastly will be the non-negative constraint and so x1+x2+x3+x4+x5+x6+x7>=0

Solving the above in excel using the solver function we get the following optimal solution:

X1 = 25
X2 = 20
X3 = 0
X4 = 0
X5 = 35
X6 = 0
X7 = 10

The minimized no. of the objective function is 90.  

The number of call center employees that exceed the minimum requirement is the number above the given minimum. So for instance the minimum number desired for Monday is 50 while x1+x4+x5+x6+x7 = 70. Thus excess employees = 70-50. Similarly we have computed the same for all other days as well:

Excess employees
Monday = 20
Tuesday = 0
Wednesday = 10
Thursday = 0
Friday = 5
Saturday = 0
Sunday = 0

Excel's solver image:

File Home Insert Page Layout q quant 2019 20 - Microsoft Excel Formulas Data Review View Connections * Clear Properties Reapp